# Latent Beam Diffusion Models for Decoding Image Sequences

### v1

@@ 3.2.3 Contrastive Score
To train the classifier effectively, we employ a contextual sequence training approach, which enables the model to capture sequential relationships across different tasks. During training, the true label $l_{t,j}$ indicates whether the candidate image $I_{j}$ fits correctly into the sequence, given the previous steps $B_{j-1}^{(b)}$ of the $b$-th beam. Here, $t$ refers to the task index and $j$ represents the step index within the sequence. The model’s prediction $\varphi(I_{j},s_{j},B_{j-1}^{(b)})$ is based on the candidate image $I_{j}$, its textual description $s_{j}$, and the context provided by the prior steps. The model aims to minimize cross-entropy loss by aligning the candidate image with its context, ensuring it fits well within the sequence according to the true labels $l_{t,j}$, defined as:
$\small\operatorname*{argmin}\sum_{t=1}^{M}\sum_{j=1}^{N}l_{t,j}\log(\varphi(I_{j},s_{j},B_{j-1}^{(b)}))$ (9)
Through contrastive training, the model learns to align representations of similar steps while enhancing the distinction between steps that differ semantically or visually.
3.2.4 Contextual Scene Descriptions
In many cases, step descriptions alone may lack context as precise visual prompts. These step descriptions are often interdependent, relying on prior steps for contextual understanding, or may miss visual details, e.g. ”Put the ice cream in freezer for 1 hour”, or actions that involve multiple tasks, such as ”Slice the tomatoes, grate the cheese, and stir the sauce”. To address this challenge, we leverage Gemini [33] to refine step descriptions into visually detailed prompts while considering sequence context. Specifically, Gemini produces a contextualized caption $c_{j}$ based on the current step description $s_{j}$ and all preceding steps as follows:
$\small c_{j}=\phi(s_{j}|\{s_{1},s_{2},\dots,s_{j-1}\})$ (10)
Following this approach we can improve image generation prompts by maintaining consistency and preserving the dependencies between sequential steps. For further information, see Appendix E.
@@ 3.2.4 Contextual Scene Descriptions
Method Human (%) Gemini (%) Human (%) Gemini (%) Human (%) Gemini (%)
Greedy/CoSeDlen=1 [32] 13.3 16.7 20.0 10.0 27.2 16.7
Nucleus Sampling [14] 33.3 16.7 30.0 30.0 18.2 26.6
GILL [22] 0.0 3.3 10.0 20.0 0.00 0.0
BeamDiffusion 53.3 63.3 40.0 40.0 54.5 56.7
Table 1: Frequency of method selection by human annotators and Gemini.
4 Experimental Setup
In this section, we describe the experimental setup to evaluate BeamDiffusion’s performance in generating coherent image sequences. We outline the datasets used, identify other baseline methods, and explain the human and Gemini evaluations focused on semantic and visual consistency across generated sequences. All data and implementations will be released after publication.
Datasets.
We used a dataset consisting of publicly available manual tasks in the Recipes and DIY domains [4], as well as the Visual Storytelling (VIST) dataset [16]. Recipes serve as the in-domain dataset, while DIY tasks and VIST are out-of-domain. These out-of-domain datasets allow us to assess the generalizability of our model across different task domains. More details can be found in Appendix D.
Baselines.
To assess the performance of our image sequence generation method for multistep tasks, we compared BeamDiffusion to two other decoding approaches, as well as GILL [22], focusing on different strategies for generating consistent image sequences:
  • Greedy/CoSeDlen=1 [32]: In this method, a new image is generated at each step using the first five latents from the previous image, selecting the best overall image. Although similar to CoSeDlen=1, it relies solely on CLIP without alterations, whereas CoSeD uses a trained module for selection.
  • Nucleus Sampling [14]: Rather than choosing the single most likely image, nucleus sampling randomly selects from the set of images whose cumulative probability exceeds a specified threshold $p$, according to CLIP. This allows for more diversity in the generated sequence while still emphasizing high-probability outcomes.
  • GILL [22]: Unlike the other methods, GILL does not explicitly search the latent space. Instead, it ensures coherent image sequences by iteratively generating images, while maintaining consistency across the sequence. GILL achieves this by deciding when to retrieve an image from a candidate set or when to generate a new image, based on the context of the previous images and text inputs.
Human Evaluation.
In the first task, the annotators compared pairs of beam search configurations in a ”battle-style” format to select which beam configuration was the best regarding beam width ($w$ in 7) and steps back ($n$ in 7). The winning configuration is compared against other configurations until a new configuration is selected as better. The selection criteria focused on semantic alignment and visual consistency. Semantic alignment assesses how well the image sequence matches the corresponding text, while visual consistency evaluates the coherence of the visuals across the sequence (e.g., keeping backgrounds, objects, or ingredients consistent).
Having chosen the optimal configuration, we need to assess how it compares to other methods. In the second annotation task, BeamDiffusion was compared with our baseline models. The annotators independently assessed the sequences, selecting the best method according to semantic and visual consistency, with the possibility of choosing multiple options or none. For further information, see Appendix C.
Gemini Evaluation.
To further validate our findings and annotate a wider range of tasks, we used Gemini for both evaluation of configurations and baselines. Its selection process closely mirrored that of human annotators, with one key difference: Gemini tends to default to ”both good” when given the option. To avoid this, we removed the ”both good” and ”both bad” options from its evaluation to ensure more accurate results.
We also refined our method for determining the best generation technique. Initially, presenting all four sequences at once caused the model to focus primarily on the last two, leading to biased results. To address this, we switched to pairwise evaluations, comparing sequences two at a time, which provided a more balanced and reliable assessment. For details, please refer to Appendix C.
Human-Gemini Agreement.
We conducted an alignment assessment between human annotators and Gemini. We asked 10 annotators to rate 16 Recipes, 10 DIYs, and 10 VIST tasks. After this assessment, we computed Fleiss’ kappa [8], obtaining a value of 70%, indicating a substantial agreement. Given this substantial alignment, we extended the evaluation to the remaining sequences with Gemini.
@@ 4 Experimental Setup
5 Results and Discussion
In the following sections, we demonstrate the performance of BeamDiffusion and examine different aspects: hyperparameters, non-linear decoding properties and a qualitative analysis.
5.1 General Results
The results in Table 1 compare the performance of four methods, Greedy/CoSeDlen=1, Nucleus Sampling, GILL, and BeamDiffusion, evaluated through human annotations and the Gemini model across three domains: Recipes (in-domain), DIY (out-of-domain), and VIST (out-of-domain).
Across domains, BeamDiffusion consistently outperforms other methods. In the Recipes domain, it leads with 53.3% in human evaluation and 63.3% in Gemini. Nucleus follows with 33.3% (human) and 16.7% (Gemini), while greedy scores 13.3% (human) and 16.7% (Gemini). In DIY, BeamDiffusion achieves 40.0% in both evaluations, again outperforming greedy and nucleus. In VIST, BeamDiffusion leads with 54.5% (human) and 56.7% (Gemini), with greedy and nucleus scoring significantly lower. GILL consistently underperforms, receiving 0% selection in VIST across both evaluations. In Recipes, it scores just 3.3% in Gemini and remains the weakest performer across all methods. While it performs slightly better in DIY, it still lags behind other approaches. Although BeamDiffusion excels in Recipes and VIST, performance in DIY remains more challenging, with smaller performance gaps between methods. This reflects the inherent complexity of the DIY domain, where generating coherent outputs is particularly difficult.
Overall, BeamDiffusion stands out as the most effective method across all domains, particularly in Recipes and VIST. However, its performance in DIY highlights the need for a more effective search in the latent space to ensure consistency and coherence across diverse tasks.
Figure 5: Example of the Recipes domain task ”Jamaican Callaloo With Shrimp”, with a sequence of 4 steps.
5.2 Best Beam Search Configuration
The hyperparameters beam width ($w$ in Eq. 7) and steps back ($n$ in Eq.  2), play a crucial role in the performance of BeamDiffusion. In this section, we evaluated different configurations across four beam widths (2, 3, 4, and 6) and three steps back (2, 3, and 4). In this study, we specifically focus on latents indexes 0, 1, 2, and 3, denoted as $K$ in Eq. 2, which yield the most promising results. A more detailed discussion on the impact of latents is provided in Appendix A.1.
We first performed human annotations on six tasks to identify the optimal beam search configuration, see Appendix C for details. The results in Table 3 showed that the best performing configuration was 2 steps back with a beam width of 4, achieving a selection rate of 25%. Other configurations, such as 3 steps back with a beam width of 2 and 4 steps back with a beam width of 2, also showed competitive performance (16.67%), but none outperformed the 2 steps back and beam width of 4 configuration. The configurations with beam width 6 performed the worst in all the settings tested.
5.3 Non-Linear Sequence Denoising
To analyze how beam diffusion explore the latent space as formulated in Sec. 3.2.1, we examined the step from which the latent representation was selected during the generation process. For an analysis of latent selection, see Appendix A.2. We focused the step selection process in four-step sequences (Figure 4). Step 1 was chosen 69.9% of the time, highlighting its key role in shaping the sequence. Selection rates decrease for later steps, with Step 3 at 46.8% and Step 2 at 33.3%. The sum of these percentages exceeds 100% because they represent the relative selection frequency of each step, considering the number of possible selection opportunities. Earlier steps, such as Step 1, have more chances of being chosen, leading to a higher selection rate. Thus, the percentages reflect how often each step is selected given the different possibilities for their inclusion, rather than a simple additive total. This pattern demonstrates how beam search prioritizes crucial steps, particularly Step 1, to establish sequence coherence while strategically weighing intermediate stages. By balancing local accuracy with global consistency, BeamDiffusion ensures high-quality image sequences that remain visually and contextually aligned.
Figure 6: Example of the out-of-domain DIY task ”How to Make a Door”, with a sequence of 5 steps.
Figure 7: Example of the out-of-domain VIST story ”July 4th Barbecue”, with a sequence of 5 steps.
@@ 5.4 Qualitative Analysis
To illustrate how different decoding methods impact the semantic and visual consistency of generated output, we present several examples in Figures 57, and in Appendix F.2. In Figure 5, we see that BeamDiffusion consistently maintains the same pan across all steps, a quality not always achieved by greedy or nucleus methods. Meanwhile, GILL struggles to follow the text from step to step, leading to more inconsistencies in the outputs.
Looking at the out-of-domain DIY and VIST datasets, in Figure 6, we observe a task with intricate details. Although other methods fail to consistently capture the color of the plywood, BeamDiffusion outperforms by maintaining this detail, demonstrating its strength in preserving fine-grained elements in complex tasks. In Figure 7, we again observe significant inconsistencies with the greedy and nucleus methods, which alternate between black-and-white and color images. In contrast, BeamDiffusion keeps the color consistent throughout the sequence. Additionally, background elements, such as the bushes, remain stable across the first three steps. Even the persons in first step closely resemble those in the third, further emphasizing the BeamDiffusion consistency. At the same time, GILL consistently struggles to follow the text, failing to capture the necessary details.
This analysis highlights BeamDiffusion ability to maintain high levels of semantic and visual consistency, even in out-of-domain tasks, much like its performance on Recipes.
@@ 6 Conclusions
Through both human and automatic evaluation, we show that beam search performs effectively in generating high-quality and coherent image sequences. The insights drawn from these evaluations reveal that beam search aligns closely with textual instructions while ensuring visual continuity. In contrast, GILL, the only method in our study that does not perform latent search, struggles to maintain consistent alignment with the text instructions, resulting in more inconsistent and less coherent image sequences. Furthermore, our results underscore the significance of the non-linear selection process in beam search, highlighting how prioritizing certain steps and latents can optimize the quality of the generation process.
Appendix A Impact of Latent Selection on Beam Search Performance

### v2

@@ 3.1.2 Pruning Diffusion Beam Paths
The contrastive classifier $\varphi(I_{j}|\{s_{1},\ldots,s_{j}\},\{I_{1},\ldots,I_{j-1}\})$ is trained as a next image prediction task to estimate how well align an image is with the previously sequence of steps and images. For a given sequence $t$ and sequence step $j$, the ground-truth label $l_{t,j}$ indicates whether image $I_{j}$ belongs to the sequence given its prompt $s_{j}$ and context $B_{j-1}^{(b)}$. Negative examples are hallucinated from other sequences. The model minimizes the cross-entropy loss encouraging the model to promote candidates that best match the sequence’s semantic and visual dynamics. See appendix A for implementation details.
4 Experimental Setup
In this section, we describe the experimental setup to evaluate BeamDiffusion’s performance in generating coherent image sequences. We outline the datasets used, identify other baseline methods, and explain the human, automatic and LLM-as-a-judge (Gemini [34]) evaluations focused on semantic and visual consistency across generated sequences.
Datasets.
We used a dataset consisting of publicly available manual tasks in the Recipes and DIY domains [2], as well as the Visual Storytelling (VIST) dataset [13]. Recipes serve as the in-domain dataset, while DIY tasks and VIST are out-of-domain. These out-of-domain datasets allow us to assess the generalizability of our model across different task domains. More details can be found in Appendix B.2.
Baselines.
To evaluate the performance of our image sequence generation method, we compared BeamDiffusion with SD2.1 [35] against two other competing approaches to generate image sequences: GILL [18] and StackedDiffusion [26]. Additionally, we considered two decoding strategies: Greedy and Nucleus Sampling [11]. As these decoding strategies are novel in the context of image generation, we provide a brief explanation of each. All methods used 50 denoising iterations.
In the Greedy approach, a new image is generated at each step using the first five latents from the previous image, with the image that scores highest overall being selected. Although this is similar to CoSeDlen=1 [33], it differs in that it relies exclusively on CLIP for selection without any learned module, whereas CoSeD incorporates a trained selection mechanism.
In contrast, Nucleus Sampling introduces diversity by sampling from the subset of images whose cumulative CLIP-assigned probability exceeds a predefined threshold $p$. Rather than always selecting the most likely image, this method favors high-probability outcomes while allowing for variation in the generated sequence.
Human Evaluation.
To find the optimal hyperparameters, the annotators compared pairs of beam search configurations to select the optimal beam width ($w$ in Eq. 5) and steps back ($m$ in Eq. 2). The winning configuration is compared against other configurations until a new configuration is selected as better. The selection criteria focused on semantic alignment and visual consistency. Semantic alignment assesses how well the image sequence matches the corresponding text, while visual consistency evaluates the coherence of the visuals across the sequence (e.g., keeping backgrounds, objects, or ingredients consistent).
Having chosen the optimal configuration, we compared BeamDiffusion to other methods. In the second annotation experiment, BeamDiffusion was compared with other baseline models. Annotators independently assessed the sequences, selecting the best method according to semantic and visual consistency with the possibility of choosing multiple options or none. For further information, see Appendix B.1.
Gemini Evaluation.
To further validate our findings and annotate a wider range of tasks, we used Gemini for both evaluation of configurations and baselines. Its selection process closely mirrored that of human annotators. For details, please refer to Appendix B.1.
Human-Gemini Agreement.
We conducted an alignment assessment to measure the agreement between human annotators and Gemini. For this, human annotators were asked to rate 16 Recipes, 10 DIYs, and 10 VIST tasks. These ratings were then compared to the evaluations made by Gemini. To quantify the agreement between the two, we computed Fleiss’ kappa [6], obtaining a value of 70%, which indicates a substantial level of agreement. Based on this strong alignment, we extended the evaluation to the remaining sequences using Gemini.
Automatic Metrics.
Following prior work [2, 33, 21], we evaluate each method using: (1) Visual Consistency, based on average image similarity via DINO [27] (DINO-I) and CLIP [31] (CLIP-I); and (2) Semantic Alignment, using CLIP (CLIP-T) to measure similarity between the current image and the average of all step texts up to and including the current one. To place the scores on a comparable scale, we normalize CLIP-I, CLIP-T and DINO by their maximum. Finally, we compute CLIP* as the product between the CLIP-I and CLIP-T, and DINO* as the product between the DINO-I and CLIP-T, this formulations balances visual and semantic quality, while emphasizing semantic alignment, which is critical for accurately reflecting the stepwise input text. See Appendix B.1 for details.
5 Results and Discussion
In this section, we describe the performance of BeamDiffusion based on both human and automatic evaluations. We begin by analyzing the human evaluation, proceed to discussing the automatic evaluations, then focus on specifics of our approach such as hyperparameters and non-linear decoding characteristics, and finally provide a qualitative analysis. Additional statistical information about the results can be found in Appendix B.1.
5.1 General Results
Table 1 compares all methods across three domains: Recipes (in-domain), DIY (out-of-domain), and VIST (out-of-domain). BeamDiffusion consistently outperforms other methods. In the context of Recipes, BeamDiffusion is selected 50.0% of the time by humans and 37.6% by the LLM, with Nucleus and Greedy following. In out-of-domain areas, it is chosen once more 50.0% of the time by humans and 30.0% by the LLM-judge for the DIY domain, and 58.3% for humans and 34.9% by the LLM for the VIST dataset. Among sequence generation methods, GILL consistently underperforms across all domains, receiving at most 3.3% of human votes and 16.1% from the LLM in VIST. StackedDiffusion performs better than both Greedy and GILL, but still lags behind BeamDiffusion, with 20.0% of human selections and 9.8% of LLM-judge selections in the Recipes domain.
Overall, BeamDiffusion excels in Recipes and VIST, however performance in DIY remains challenging, with smaller gaps between methods due to the domain’s inherent complexity.
Table 1: Frequency of method selection by human annotators(H) and Gemini(G).
@@ 5.1 General Results
Method H (%) G (%) H (%) G (%) H (%) G (%)
GILL [18] 0.0 2.0 0.0 20.0 3.3 16.1
StackedDiffusion [26] 20.0 9.8 12.5 15.6 0.00 3.2
Image Sequence Decoding
Greedy/CoSeDlen=1 [33] 0 27.6 0 23.3 0 26.3
Nucleus Sampling [11] 30.0 23.0 37.5 27.7 8.3 26.6
BeamDiffusion 50.0 37.6 50.0 30.0 58.3 34.9
5.2 Automatic Metrics
Table 2 shows that BeamDiffusion achieves the highest CLIP* and DINO* scores (0.40 and and 0.30), demonstrating a strong balance between image consistency and semantic alignment. BeamDiffusion excels in maintaining a good compromise between text grounding and image similarity achieving the best result in image consistency (CLIP-I of 0.83 and DINO-I of 0.62). Similarly, for CLIP-T, BeamDiffusion achieves the highest score for image-text alignment (CLIP-T of 0.48). However, in this case, the highest result is shared, as Greedy and Nucleus obtain the same score. This outcome is expected because Greedy and Nucleus utilize CLIP-T to select their outputs by aligning the generated sequence with the text. In terms of sequence generation methods, StackedDiffusion performs better than GILL in all metrics but still lags behind BeamDiffusion and the other decoding methods.
Table 2: Comparison of CLIP-I, DINO-I and CLIP-T across different methods
Method CLIP-I DINO-I CLIP-T CLIP* DINO*
GILL 0.55 0.49 0.43 0.23 0.21
StackedDiffusion 0.76 0.46 0.47 0.36 0.22
Image Sequence Decoding
Greedy 0.81 0.54 0.48 0.39 0.26
Nucleus 0.77 0.53 0.48 0.37 0.25
BeamDiffusion 0.83 0.62 0.48 0.40 0.30
5.3 Best Beam Search Configuration
The hyperparameters beam width ($w$ in Eq. 5) and steps back ($m$ in Eq. 2), play a crucial role in the performance of BeamDiffusion. In this section, we report an exaustive evaluation on how different beam widths (2, 3, 4, and 6) and steps back (2, 3, and 4) improve the latent selection and the overall quality of the generation. A more detailed discussion on the impact of the latents selection is provided in Appendix C.
We first perform human annotations on a set of sequences to identify the optimal beam search configuration, see Appendix B.1 for details. To determine the best configuration, annotations were used to discard non-optimal configurations. The results in Table 4 showed that the best performing configuration was 2 steps back with a beam width of 4, achieving a selection rate of 25%. Other configurations, such as 3 steps back with a beam width of 2 and 4 steps back with a beam width of 2, also showed a strong performance (16.67%), but none outperformed the 2 steps back and beam width of 4 configuration. The configurations with beam width 6 performed the worst in all the settings tested.
@@ 5.3 Best Beam Search Configuration
5.4 Non-Linear Sequence Denoising
To analyze how BeamDiffusion explores the latent space, we measured the number of times that a sequence step provides latents for the generation of subsequent steps of the sequence. In this experiment we consider sequences of four steps, (Figure 2). Appendix C.2 provides an analysis at the level of the denoising iterations. Latents from step 1 were chosen 69.9% of
@@ 5.4 Non-Linear Sequence Denoising
the times they had a chance of being selected, highlighting the role of step 1 in shaping the sequence. Selection rates decrease for later steps, with Step 3 at 46.8% and Step 2 at 33.3%. Earlier steps, such as Step 1, have more chances of being chosen, leading to a higher selection rate. Thus, the percentages reflect how often each step is selected given the different possibilities for their inclusion, rather than a simple additive total. This pattern demonstrates how beam search prioritizes crucial steps, particularly Step 1, to establish sequence coherence while strategically weighing intermediate sequence steps.
(a) Recipes domain task.
(b) DIY task.
(c) VIST story.
Figure 3: Example tasks from three different domains: Recipes ("Jamaican Callaloo With Shrimp"), DIY:("How to Make a Door"), and VIST ("July 4th Barbecue").
@@ 5.5 Qualitative Analysis
To highlight how different decoding methods and models impact the semantic and visual consistency of generated outputs, we present several examples in Figures 3(a)3(c) and Appendix F.2. We compare three decoding methods (Greedy, Nucleus, and Beam) with two models (GILL and StackedDiffusion).
In Figure 3(a), BeamDiffusion consistently maintains the same pan across all steps, a feature not always seen in Greedy or Nucleus methods. While GILL struggles to follow the step text and results in more inconsistencies, Beam performs better in preserving visual coherence compared to Greedy and Nucleus, though it still faces some challenges. For DIY tasks,as shown in Figure 3(b), BeamDiffusion can preserve basic elements such as the color of the plywood and a basic adherence to the text. However, overall, all methods face challenges in capturing precise details. In the VIST dataset, Figure 3(c), Greedy and Nucleus alternate between black-and-white and color images, while BeamDiffusion ensures color consistency throughout the sequence. Furthermore, background elements, such as bushes, are maintained in the first three steps, and the individuals in the first and third steps closely resemble each other. Overall, BeamDiffusion excels in maintaining both semantic and visual consistency, even in complex out-of-domain tasks, In contrast, methods such as StackedDiffusion and GILL encounter challenges in keeping a balance between text alignment and consistency across images.
6 Limitations
Exploring the latent denoising space with BeamDiffusion requires Diffusion Models that provide access to fine-grain latents across denoising iterations, where BeamDiffusion still has the opportunity to steer the reverse diffusion process.
The theoretical complexity of BeamDiffusion is linear with the number of beams that we configure the model to explore. The drawback of this strategy is that exploring the space of hypothesis increases the computational complexity of the process. However, this provides a solid ground to explore exciting new ideas where stronger constraints can be enforced in the explored beams.
@@ 7 Conclusions
Through both human and automatic evaluation, we show that beam search performs effectively in generating high-quality and coherent image sequences. The insights drawn from these evaluations reveal that beam search aligns closely with textual instructions while ensuring visual continuity. Most methods can align images with text, however, we demonstrated that sequence decoding methods perform naturally better at aligning their output with the input sequence, and BeamDiffusion in particular achieves an optimal balance between the sequence of text instructions and the visual coherence of the generated images. Furthermore, our results underscore the significance of the non-linear selection process in beam search, highlighting how prioritizing certain steps and latents can optimize the quality of the generation process.
This work shows that the latent diffusion space can be steered, offering more reliable and factual outputs. As future work, we plan to explore constraint decoding method that assess the likelihood of each denoise latent with respect to the target prompt. Solving this challenge, will improve the correlation between all elements in a sequence of scenes.
Appendix A Implementation Details
Base Model.
Our method is implemented on top of Stable Diffusion v2.1 [35], a latent diffusion model for high-quality text-to-image synthesis. We build upon the open-source implementation provided by Stability AI, extending it to support beam-based generation, latent trajectory tracking, and contrastive scoring mechanisms as described in the main text.
Latent Trajectory Reuse.
In our approach, maintaining coherence across sequential generations requires access to the intermediate latent representations produced during the generative process. Fortunately, the model pipeline inherently provides the full sequence of latent representations at each generation step. This enables us to easily extract and reuse specific latents as seed for new generations.
Classifier Training.
To train the contextual classifier, we employ a sequence-aware training approach that allows the model to capture dependencies between steps in the generation process. The classifier $\varphi$ takes as input a candidate image $I_{l}$, its textual description $s_{l}$, and the previous sequence images and the respective step text ($B_{l-1}^{(b)}$). Then the classifier scores $I_{j}$ based on how well it fits within the evolving trajectory. During training, we use labels $l_{d,l}$ indicating whether $I_{l}$ is the correct continuation in the sequence for task $d$ and step $l$. The model is trained to minimize the cross-entropy loss:
$\operatorname*{argmin}\sum_{d=1}^{D}\sum_{l=1}^{L}l_{d,l}\cdot\log(\varphi(I_{l},s_{l},B^{(b)}_{l-1})),$ (6)
This contrastive training encourages the model to assign high scores to contextually appropriate continuations while penalizing incoherent or semantically unrelated ones. As a result, the classifier learns to distinguish subtle differences in step-to-step coherence, both visually and textually.
Hardware.
All experiments were ran on NVIDIA A100 GPUs with 40GB of memory.
Appendix B Experiments Details
B.1 Metrics and Annotations
Automatic Metrics.
In image sequence generation, CLIP similarity scores between visually similar frames often exceed 0.85, even for near-duplicate content. Wang et al. [49] highlight that such high CLIP-I scores can indicate semantic duplication, which poses a challenge when diversity across frames is desired. While CLIP-I and DINO-I are common proxies for semantic similarity, rewarding near-duplicates undermines meaningful visual progression. Conversely, poor alignment between images and their textual prompts is reflected in low CLIP-T scores. Lin et al. [20] address this by clipping CLIP-T values below 0.1, effectively filtering out weak text-image matches.
Figure 4: Illustration of the limitation of CLIP-I: despite a high CLIP-I score, the result is not semantically correct or visually faithful. The example compares several automatic metrics (CLIP-T, CLIP-I, DINO-I, CLIP*, DINO*), highlighting discrepancies and failure cases in sequence evaluation.
To balance these factors, we set CLIP-I scores above 0.9 and DINO-I scores above 0.85 to zero, penalizing visual redundancy. Similarly, we zero out CLIP-T scores below 0.1 to discourage poor alignment with the prompt. This clipping strategy encourages the generation of sequences that are both semantically relevant and visually diverse. In Figure 4, we present the automatic metric scores for a sequence across all methods. The results illustrate that very high CLIP-I values can lead to sequences that fail to follow the textual steps accurately, as observed in the case of GILL. Despite achieving strong visual similarity, the sequence diverges from the intended narrative, highlighting a key limitation of relying solely on CLIP-I . This example reinforces the importance of our clipping strategy: without penalizing excessively high CLIP-I values, the generation process may favor near-duplicates over meaningful visual progression, ultimately undermining semantic fidelity to the prompt.
Table 5: Comparison of CLIP-I, DINO-I and CLIP-T across different methods with the standard deviation
Method CLIP-I DINO-I CLIP-T
GILL 0.55 $\pm$ 0.30 0.49 $\pm$ 0.26 0.43 $\pm$ 0.08
StackedDiffusion 0.76 $\pm$ 0.12 0.46 $\pm$ 0.14 0.47 $\pm$ 0.09
Image Sequence Decoding
Greedy 0.81 $\pm$ 0.13 0.54 $\pm$ 0.14 0.48 $\pm$ 0.07
Nucleus 0.77 $\pm$ 0.15 0.53 $\pm$ 0.15 0.48 $\pm$ 0.08
BeamDiffusion 0.83 $\pm$ 0.14 0.62 $\pm$ 0.18 0.48 $\pm$ 0.09
To complement average scores, we analyze the standard deviation ($\sigma$) of these metrics (Table 5). A low $\sigma$ reflects stable, consistent performance between diverse sequences, while a higher $\sigma$ implies variability in quality or adaptation to task complexity.
GILL exhibits the highest $\sigma$ in CLIP-I and DINO-I, indicating its semantic similarity varies significantly between tasks, likely due to inconsistent adherence to step instructions. StackedDiffusion’s low $\sigma$ indicates its consistent performance across different tasks. BeamDiffusion strikes a balance with moderate $\sigma$ values, indicating a balance between consistency and variation across tasks.
It is worth noting that CLIP-T generally shows lower standard deviation across models, since this metric is explicitly minimized during the diffusion model training. As a result, CLIP-T scores tend to be more stable regardless of the task-specific variation.
Together, the mean and standard deviation metrics suggest BeamDiffusion offers the best combination of high semantic fidelity, text alignment, and stable adaptation across different sequences, making it the most robust method for sequential image generation.
Human annotations.
  • Steps Back ($m$ in Eq. 2): This refers to how many previous steps the model uses from its latent representations to explore new candidate outputs. We evaluated three different settings for the number of steps back: 2, 3, and 4, which correspond to using latents from the last two, three, or four images, respectively.
  • Latents Indexes: We tested two sets of latent configurations, 1, 3, 5, 7 and 0, 1, 2, 3, by selecting specific indexes from the latent representations generated in previous stages. This was done to explore how varying levels of latent detail impact the quality of the generated sequences.
  • Beam Width ($w$ in Eq. 5): We explored four different beam widths, 2, 3, 4 and 6, which determine how many beams are retained at each step during the search process.
Evaluating all of these at once proved to be exhausting and difficult to manage. To simplify the evaluation process, we adopted a elimination annotation method. In this approach, 5 annotators compare two configurations at a time and select the one with better overall coherence. If neither configuration stands out, annotators can select "both good" or "both bad" (Figure 5), in which case the computationally cheaper option advances to the next round. This process is repeated until we identify the best configuration. This method focuses on direct comparisons, making it easier to pinpoint the most effective beam search settings while also considering computational efficiency.
Figure 5: Example of the human annotation elimination round for selecting the best beam search configuration.
To evaluate all models performance, annotators are presented with five generated sequences side by side. Figure 6 illustrates an example of this comparison process, where the annotators assess the sequences to determine the most effective approach.
Figure 6: Example of a human annotation to choose the best method.
Gemini evaluation.
Gemini’s selection process for annotating the best beam search configuration and the best generation method closely mirrored that of human annotators, with one key adjustment: Gemini tends to default to "both good" or "both bad" when such options are available. To ensure more decisive and accurate results, we removed these options from the prompt. We also refined our comparison methodology. Initially, presenting all five sequences at once led to biased evaluations, as the model disproportionately favored the last two sequences. To mitigate this, we adopted a pairwise evaluation strategy, comparing two sequences at a time. This approach resulted in more balanced and reliable assessments across all comparisons.
Figure 7 shows the prompt used in Gemini to guide the evaluation of image sequences based on visual and semantic consistency.
Gemini Prompt: "Evaluate two image sequences using the following criteria: Visual Consistency: Objects, ingredients, backgrounds, colors and styles should remain consistent across frames. Frames should transition smoothly, resembling a video-like progression with no sudden changes or missing elements. Semantic Consistency: Images must logically follow the described steps, clearly and accurately. Steps of the process: {steps} After evaluating all sequences, select the one with the strongest overall consistency (visual and semantic). Pay close attention to even minimal differences, as the sequences are very similar. If both sequences are equally good or bad, choose the one that is slightly better. Choose exactly one of the following options: - "First sequence" - "Second sequence" "
B.2 Datasets
Recipes and DIYs.
VIST.
Appendix C Impact of Latent Selection on Beam Search Performance